OpenStudy (anonymous):

solve sqrt{7x+2}= 2x and then check answers.

OpenStudy (anonymous):

As with the last one, square both sides (and remember that when you do so, you have to choose the answers carefully at the end): \[\sqrt{7X+2} = 2x\] \[7X + 2 = 4X^{2}\] \[4X^{2} - 7X - 2 = 0\]

OpenStudy (anonymous):

\[(\sqrt{7x+2})^2=(2x)^2\]

OpenStudy (anonymous):

this one gives a solution, unlike last one

OpenStudy (anonymous):

You'll note that when you solve the quadratic equation you'll get x=2, x=-1/4; that's because we squared both sides. But you can't have the square root of a negative number, so disregard x=-1/4

OpenStudy (anonymous):

elai will work it out

OpenStudy (anonymous):

oh because \[4x^2-7x-2=(x-2)(4x+1)=0\]

OpenStudy (anonymous):

so x = 2 is your answer but you have to check it in original equation \[\sqrt{7\times 2+2}=2\times 2\] \[\sqrt{16}=4\] \[4=4\]

OpenStudy (anonymous):

so i only get one answer of 2 right? nd it turns out to be right?(:

OpenStudy (anonymous):

yes one answer only and it is x = 2